OpenStudy (anonymous):

Find a cubic function f(x)=ax^3+cx^2+d that has a local maximum value of 10 at -4 and a local minimum value of 7 at 0

OpenStudy (danjs):

you are given 2 points on the function,

OpenStudy (anonymous):

ya I realize that

OpenStudy (anonymous):

I just don't know how to plug them in to solve since there are several variables

OpenStudy (danjs):

which class

OpenStudy (danjs):

put in both points, see what you get first in terms of a system of equations in the constants a,c,d

OpenStudy (danjs):

10 = -64a + 16c + d d = 7

OpenStudy (danjs):

is this calc class, or before derivatives

OpenStudy (danjs):

because you have 2 variables but one equation, they give you max and min points though, the derivative is 0 at those

OpenStudy (anonymous):

I'm still kind of lost

OpenStudy (danjs):

use the two points given (x , f(x)) and you get an equation for each one 10 = -64a + 16c + d 7 = d

OpenStudy (danjs):

overall, you have one equation and 2 variables left, cant solve , need another relationship with a and b ,

OpenStudy (danjs):

the derivative is f ' (x ) = 3a*x^2 + 2c*x Using this because they tell you those points are max/min locations, here f ' (x) = 0

OpenStudy (danjs):

So at the point (-4 , 10) there is a horizontal tangent line, the first derivative is zero, it is a maximum value you can use x=-4, and f ' (x)=0, that will give you another relationship between 'a' and 'b', might be able to get a solution now

OpenStudy (danjs):

the two equations you have now are 10 = -64a + 16c + 7 ---(d=7) and 0 = 48a - 8c may get values for 'a' and 'b' after solving those

OpenStudy (danjs):

a and c , sorry

OpenStudy (danjs):

i got a=3/32 and c=9/16, and d =7 from before

OpenStudy (danjs):

you follow all that?

OpenStudy (anonymous):

thanks!